OpenStudy (anonymous):

find dy/dx of sin(x+y) + sin (x-y) = 1 a. -cotxcoty b. cotxcoty c. 1 d. -1

myininaya (myininaya):

\[(x+y)_xcos(x+y)+(x-y)_x \cos(x-y)=1_x\] \[(x_x+y_x)\cos(x+y)+(x_x-y_x)\cos(x-y)=0\] \[(1+y')\cos(x+y)+(1-y')\cos(x-y)=0\] Distribute \[\cos(x+y)+y' \cos(x+y)+\cos(x-y)-y'\cos(x-y)=0\] \[y'(\cos(x+y)-\cos(x-y))=-\cos(x-y)-\cos(x+y)\] do you think you can finish this?

OpenStudy (anonymous):

-1?

myininaya (myininaya):

\[\frac{-(\cos(x-y)+\cos(x+y))}{\cos(x+y)-\cos(x-y)}\]

myininaya (myininaya):

so you have to do some trig identity stuff

OpenStudy (anonymous):

(-cos(x-y)-cos(x+y))/(y-cos(x)+cos(x+y)) then after this?

myininaya (myininaya):

\[=-\frac{\cos(x) \cos(y)+\sin(x) \sin(y)+\cos(x) \cos(y)- \sin(x) \sin(y)}{\cos(x) \cos(y) - \sin(x) \sin(y)-\cos(x) \cos(y) - \sin(x) \sin(y)}\]

myininaya (myininaya):

\[=-\frac{2 \cos(x) \cos(y)}{-2 \sin(x) \sin(y)}\]

myininaya (myininaya):

\[=\frac{\cos(x)}{\sin(x)} \cdot \frac{\cos(y)}{\sin(y)}\]

myininaya (myininaya):

got it from here now?

OpenStudy (anonymous):

thats y u said i need trigo identity ic ic from my trigo identity on my note book it is cotxcoty

myininaya (myininaya):

lol yep :)

myininaya (myininaya):

hey phi can yo help i must leave sorry guys

OpenStudy (anonymous):

thank myininaya

OpenStudy (phi):

y = (sin2x)/4 + x/2 find dy/dx

OpenStudy (anonymous):

x = pi /2 ryt?

OpenStudy (phi):

that's what I mean find dy/dx for y= 1/4 sin(2x) +x/2 dy/dx = (1/4) * cos(2x) * 2 *dx/dx + 1/2 dx/dx = cos(2x)/2 + 1/2 now solve the equation cos(2x)/2 + 1/2 = 1 or cos(2x)+1 = 2 cos(2x) = 1 2x = acos(1) = 0 or 2 pi x= 0/2 = 0 or 2 pi/2 = pi so x= n*pi for n= an integer

OpenStudy (phi):

the only one that matches is (b) 0

OpenStudy (anonymous):

ow 6 people i thought like this also cos(2x)=-1 2x=cos−1(−1) 2x=π x=π/2

OpenStudy (anonymous):

i go for 0 now =)

OpenStudy (phi):

when in doubt, use wolfram http://www.wolframalpha.com/input/?i=d%2Fdx+%28sin2x%29%2F4+%2B+x%2F2+%3D+1

OpenStudy (anonymous):

thanks again phi